# AI in Supply Chain Risk Assessment: A Systematic Literature Review and Bibliometric Analysis

### v2

@@ 4.10 Bibliographic Coupling
Bibliographic coupling is a method in bibliometrics that assesses the similarity between documents based on their shared references. This approach identifies connections and relationships among scholarly works by analyzing the overlap in their citation patterns. In other words, if two documents cite a similar set of references, they are considered to be bibliographically coupled. This analysis helps reveal thematic associations, collaborative efforts among authors, and the intellectual structure of a research field. Clusters formed through bibliographic coupling indicate groups of documents, authors, countries, organizations, or sources that share common references, suggesting a degree of interrelatedness in their research content. The strength and nature of these couplings provide valuable insights into the structure and dynamics of academic knowledge networks.
@@ 4.10.1 Authors
In Figure 9, we identified 39 significant authors out of 1407, forming six distinct clusters based on a minimum document threshold of 3 per author. These clusters represent cohesive groups of authors who share common research interests, as evidenced by the co-citation of their works. Here influential contributors like ”Gupta, Shivam,” ”Ivanov, Dmitry,” ”Modgil, Sachin,” ”Qu, Yingchi,” ”Wang, Gang-Jin,” ”Xie, Chi,” and ”Zhu, You” were identified, signifying their substantial impact within cohesive clusters of authors, emphasizing shared research interests through their co-cited works.
Figure 9: Bibliographic coupling analysis of authors.
@@ 4.10.2 Countries
Figure 10 revealed 39 nations out of 67 that met the criteria, forming 6 clusters based on a minimum document threshold of 3 per country. These clusters provide insights into international collaborations and thematic concentrations among countries in scholarly publications. Standing out prominently are the People’s Republic of China and major contributors like England, Germany, and the USA, who demonstrate extensive involvement and collaborations in this domain, highlighting significant thematic concentrations among nations in scholarly publications.
Figure 10: Bibliographic coupling analysis of countries.
@@ 4.10.3 Documents
In Figure 11, 109 significant papers out of 434 met the criteria, forming 8 clusters based on a minimum citation threshold of 16 per document. These clusters highlight cohesive groups of highly cited papers, indicating thematic similarities and influential works. Here, influential authors with significant citation impact include ”Nayal,” ”Baryannis,” ”Ivanov,” ”Hosseini,” ”Modgil,” ”Zouari,” ”Wu,” and ”Hosseini,” suggesting their substantial contributions.
Figure 11: Bibliographic coupling analysis of documents.
@@ 4.10.4 Organizations
In Figure 12, 34 significant entities out of 769 met the dual criteria of a minimum document threshold of 5 and a minimum citation threshold of 10, forming 3 clusters. These clusters illustrate collaborative research efforts and the impact of organizations in the academic landscape. It shows institutions like ”Hong Kong Polytechnic University,” ”Neoma Business School,” and ”Indian Institute of Technology Delhi” as notable contributors to these cohesive groups.
Figure 12: Bibliographic coupling analysis of organizations.
@@ 4.10.5 Journals
For the bibliographic coupling analysis of sources, 42 significant publication sources out of 196 met the criteria, forming clusters based on a minimum document threshold of 3 per source, as shown in Figure 13. These clusters reveal patterns in scholarly publishing, highlighting sources that frequently publish related content. Journals such as ”Journal of Computational and Applied Mathematics,” ”IEEE Transactions on Engineering Management,” ”Annals of Operations Research,” and ”International Journal of Production Research” frequently publish related content within their respective clusters.
Figure 13: Bibliographic coupling analysis of publications sources.

### v3

@@ 4.10 Bibliographic Coupling
Bibliographic coupling assesses the similarity of documents, authors, or organizations based on shared references, revealing thematic alignments and intellectual structures. Key insights are as follows:
@@ 4.10.1 Authors
Figure 9 illustrates the bibliographic coupling analysis of 28 authors, grouping them into six thematic clusters based on shared references, highlighting intellectual connections in SCR research. Prominent contributors like Ivanov D, Gupta S, and Wang G-J demonstrate centrality, reflecting their significant influence in areas such as risk prediction, disruption management, and quantitative modeling. The clusters reveal interdisciplinary themes, including resilience strategies, industry-specific applications, and the integration of AI-driven methods. This analysis showcases both established research directions and emerging trends, providing a roadmap for future studies and collaborations in SCR.
Figure 9: Bibliographic coupling analysis of 28 authors.
@@ 4.10.2 Countries
Figure 10 illustrates the bibliographic coupling among 18 countries, categorizing them into clusters highlighting global research collaborations in AI-driven SCRA. Countries such as China, the United States, and Germany emerge as central hubs, showcasing their pivotal roles in advancing research and promoting international partnerships. The clusters signify thematic alignments across regions, with nations collaborating on critical topics like predictive analytics, sustainability, and resilience strategies. This visualization underscores the significance of cross-border collaboration in addressing global SC challenges through AI innovation.
Figure 10: Bibliographic coupling analysis of 18 countries.
@@ 4.10.3 Documents
Figure 11 illustrates the bibliographic coupling analysis of 62 documents, visually depicting the thematic clustering and citation relationships among the studies. The eight clusters highlight distinct research themes within the field, showcasing the interconnectedness of key works. Highly influential documents, such as those by “Baryannis G.” and “Ivanov D.,” emerge as central nodes, reflecting their significant contributions to the discourse on SCRA using AI. This visualization underscores the collaborative and evolving nature of the research domain, providing insights into how various studies align and contribute to specific thematic advancements.
Figure 11: Bibliographic coupling analysis of 62 documents.
@@ 4.10.4 Organizations
Figure 12 presents the bibliographic coupling analysis of 24 organizations, highlighting their collaborative and thematic alignments in AI-based SCRA research. The three clusters represent distinct institutional networks, with prominent contributors like the Hong Kong Polytechnic University and the Indian Institute of Technology Delhi playing central roles. This visualization emphasizes the global distribution of research efforts and showcases the leadership of key institutions in advancing AI-driven methodologies for SCRA.
Figure 12: Bibliographic coupling analysis of 24 organizations.
@@ 4.10.5 Journals
Figure 13 illustrates the bibliographic coupling analysis of the core academic journals that are driving the advancements in SCRA research by showing the interconnections and thematic clustering among 17 influential journals. Journals such as the “International Journal of Production Research” and “Annals of Operations Research” play a crucial role, providing foundational studies and methodologies that inform AI-driven approaches to SCRA. These clusters reveal thematic trends and emerging focus areas and facilitate knowledge dissemination by connecting researchers across disciplines.
Figure 13: Bibliographic coupling analysis of 17 publications sources.
Table 4: Publication Trends by Journal (2015-2024)
{tblr}
hline1-2,22 = -, Source Title & 2016 2017 2018 2019 2020 2021 2022 2023 Total
ANNALS OF OPERATIONS RESEARCH           2 4 1 7
COMPUTERS INDUSTRIAL ENGINEERING       1   1 2 2 6
INTERNATIONAL JOURNAL OF PRODUCTION RESEARCH       1     4 1 6
SUSTAINABILITY       1 1 1 1 1 5
WIRELESS COMMUNICATIONS MOBILE COMPUTING           1 3   4
EXPERT SYSTEMS WITH APPLICATIONS 1           1 1 3
INDUSTRIAL MANAGEMENT DATA SYSTEMS     1     1 1   3
INTERNATIONAL JOURNAL OF LOGISTICS MANAGEMENT           1 2   3
INTERNATIONAL JOURNAL OF PRODUCTION ECONOMICS       1 1     1 3
NEURAL COMPUTING APPLICATIONS   1       1 1   3
COMPUTATIONAL INTELLIGENCE AND NEUROSCIENCE             2   2
ENTERPRISE INFORMATION SYSTEMS         2       2
FUTURE GENERATION COMPUTER SYSTEMS       1     1   2
INTERNATIONAL JOURNAL OF DISTRIBUTED SENSOR NETWORKS       1 1       2
INTERNATIONAL JOURNAL OF OPERATIONS PRODUCTION MANAGEMENT             1 1 2
JOURNAL OF INTELLIGENT FUZZY SYSTEMS           2     2
KYBERNETES               2 2
OPERATIONS MANAGEMENT RESEARCH             2   2
ALEXANDRIA ENGINEERING JOURNAL             1   1
APPLIED SCIENCES-BASEL           1     1